# Distilling the Implicit Multi-Branch Structure in LLMs’ Reasoning via Reinforcement Learning

### v3

@@ 3.1.2 Structured Reward Mechanism
After above generation, we denote $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$ as the sequences generated from reasoning paths of the teacher LLM and the student LLM respectively:
$\mathbf{S^{t}}=[(\mathcal{M}{{}^{t}_{1}},\mathcal{Q}{{}^{t}_{1}}\&\mathcal{A}{{}^{t}_{1}}),(\mathcal{M}{{}^{t}_{2}},\mathcal{Q}{{}^{t}_{2}}\&\mathcal{A}{{}^{t}_{2}}),(\mathcal{M}{{}^{t}_{3}},\mathcal{Q}{{}^{t}_{3}}\&\mathcal{A}{{}^{t}_{3}}),...,(\mathcal{M}{{}^{t}_{n}},\mathcal{Q}{{}^{t}_{n}}\&\mathcal{A}{{}^{t}_{n}})],$
$\mathbf{S^{s}}=[(\mathcal{M}{{}^{s}_{1}},\mathcal{Q}{{}^{s}_{1}}\&\mathcal{A}{{}^{s}_{1}}),(\mathcal{M}{{}^{s}_{2}},\mathcal{Q}{{}^{s}_{2}}\&\mathcal{A}{{}^{s}_{2}}),(\mathcal{M}{{}^{s}_{3}},\mathcal{Q}{{}^{s}_{3}}\&\mathcal{A}{{}^{s}_{3}}),...,(\mathcal{M}{{}^{s}_{m}},\mathcal{Q}{{}^{s}_{m}}\&\mathcal{A}{{}^{s}_{m}})].$
We propose the Structured Reward Mechanism to map the generated sequence consisting of multiple meta-reasoning and solving to a reward value according to the alignment between $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$. Although this computation is carried out based on a linear sequence, when combined with RL, it becomes capable of quantifying the degree of alignment between the teacher and the student in terms of the implicit multi-branch structure, serving as environmental feedback in RL. This is because it can assess the step-level alignment between the sequences and, through the mechanism of reward for RL, finely guide the student LLM to make correct sampling from multiple potential sub-problems at each meta-reasoning step. Compared with SFT-based distillation, it makes the student LLM learn the implicit multi-branch structure in the authentic reasoning of the teacher LLM rather than just focuse on memorizing the reasoning path on the surface generated by the teacher LLM. Compared with existing RL methods, it can guide the student LLM on how to do better sampling at each step and avoid reward hacking Amodei et al. (2016); Di Langosco et al. (2022) by step-to-step comparison and early-exist mechanism.
Specifically, our structured reward mechanism sequentially compares the corresponding steps of $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$ (Algorithm 1). We use Qwen-2.5-7B-Instruct to determine whether two texts are matched. For the steps where $\mathcal{M}{{}^{t}_{i}}$ matches $\mathcal{M}{{}^{s}_{i}}$, we assign a temporary reward value of 111. We further deduct the temporary reward by judging the matching relationship of $\mathcal{Q}_{i}$ and $\mathcal{A}_{i}$. The temporary reward value is added to the total reward value after this. When meeting the mismatched $\mathcal{M}_{i}$, this reward accumulation process is ended. This reward mechanism follows the sequential dependency of the reasoning path and can give an approximately unique reward value for each step.
@@ 3.2 RL-based Knowledge Distillation Training
We combine our Generative Structure Reward Models (GSRM) with Group-based Relative Policy Optimization (GRPO) Shao et al. (2024) for RL-based knowledge distillation training. GRPO introduces a group-constrained policy update method to address the exploration-stability trade-off. GRPO enables differentiated policy optimization: each group applies an adaptive clipping threshold $\epsilon_{k}\propto 1/\sigma_{k}$ (where $\sigma_{k}$ is the group’s advantage standard deviation) to its policy ratio $r_{k}(\theta)$, constraining updates with:
$\mathcal{L}^{\text{GRPO}}(\theta)=\frac{1}{K}\sum_{k=1}^{K}\left[\min\left(r_{k}\hat{A}_{k},\text{clip}\left(r_{k},1-\epsilon_{k},1+\epsilon_{k}\right)\hat{A}_{k}\right)\right]-\beta\cdot\left[D_{\text{KL}}\left(\pi_{\theta}\|\pi_{\text{ref}}\right)\right],$ (2)
where $\hat{A}_{k}=A_{k}/\sigma_{k}$ is the group-normalized advantage, $K$ is the number of responses in a group. The dual-constraint design combines low-variance intra-group updates (via adaptive clipping) with global policy coherence (via KL divergence). In training, we combine the reward obtained from GSRM and the outcome reward of the specific task, such as the accuracy in math, in a weighted manner as the total reward for GRPO. Ablation study about the weights can be found in Appendix.
@@ 4 Experiments
This section conducts a series of experiments to analyze our method.
@@
We use OpenR1-Math [^1] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k as the training dataset for RL, in which Deepseek-R1 generated responses in this dataset are used as the teacher LLM’s reasoning paths. We keep the training datasets consistent with baselines including PPO and GRPO. In evaluation, we use the popular and challenging datasets on LLM’s reasoning including AIME24 MAA and MATH-500 Hendrycks et al. (2021) for math reasoning and GPQA-Diamond Rein et al. (2024) for graduate-Level Q&A. As for the metrics, we follow existing studies on LLM’s reasoning Guo et al. (2025); Shao et al. (2024); Zhang et al. (2025); Deng et al. (2024); Xu et al. (2024b) to use pass@k Chen et al. (2021). Our validation is divided into two parts: pass@1 and pass@k ($k>1$). As for pass@1, we generate $m$ ($m$ is 646464 for AIME, 888 for GPQA and 444 for MATH500) responses for each question and compute pass@1 as $pass@1=\frac{1}{m}\sum_{i=1}^{m}p_{i}$, in which $p_{i}$ is the correctness of the $i$-th response. This can alleviate randomness on small datasets. As for pass@k ($k>1$), it involves the LLM generating $k$ responses for each question, with the data sample being marked as accurate if at least one of the $k$ responses is accurate. Paying attention to the metric where $k>1$ is crucial, as in this setting, the LLM has the opportunity to explore multiple diverse paths to answer the question, it reflects the LLM’s ability to sample from multiple implicit paths during reasoning, thereby assessing whether the distilled LLM has learned authentic reasoning or merely memorized the teacher’s paths.
@@
We categorize the baselines into three groups according to the experimental settings. The first setting aims to show the effect of our method on improving LLM’s reasoning ability. We compare our RL-only training method with the LLM trained in SFT-RL pipeline. We use Qwen2.5-Math-7B-Instruct Yang et al. (2024), a powerful reasoning LLM trained on large-scale chain-of-thought datasets with SFT and then GRPO. Both our method and Qwen2.5-Math-7B-Instruct use the same based model: Qwen2.5-Math-7B, which has been pre-trained on math corpus. The second setting aims to explore whether our RL-based distillation can further improve the performance of the SFT-distilled LLM and achieve effective optimization of the SFT-RL pipeline in knowledge distillation. We compare our method with RL baselines including PPO and GRPO based on DeepSeek-R1-Distill-Qwen-7B Guo et al. (2025), a powerful LLM that is SFT-distilled from Deepseek-R1. The third setting aims to compare which method is better for the student LLM to learn authentic reasoning rather than memorizing the teacher’s path. We use embedding222obtained by gte-Qwen2-7B-instruct similarity Deng et al. (2025) to select the 3.2K data samples with the largest difference from the test set in OpenR1-Math-220k training set, and train Qwen2.5-Math-7B on this subset with RL-based distillation and SFT-based distillation respectively.
@@
In training, we build our code based on Open-R1333https://github.com/huggingface/open-r1, an open-source project for LLM’s reasoning. We use Pytorch 2.5.1 as the training framework and Deepspeed 0.15.4 for acceleration of parallel computing. In RL training, we use one Ascend 910B 64G NPU for online inference and four Ascend 910B 64G NPUs for training under deepspeed-zero3 setting. As for hyperparameters, we set per-device batch size as 222, gradient accumulation steps as 444, group size for GRPO as 444, temperature in online inference as 0.70.70.7. In evaluation, we use Lighteval [^4] https://github.com/huggingface/lighteval as the toolkit and follow settings in Deepseek Guo et al. (2025) to set temperature as 0.60.60.6, max new tokens as 327683276832768 and top-p as 0.950.950.95. We report the results of multiple runs to reduce the randomness.
(a) Performance on AIME24
@@
(b) Performance on GPQA-Diamond and MATH-500
@@
means traing set of MATH-500 has appeared in the SFT training data of Qwen2.5-Math-7B-Instruct Yang et al. (2024).
@@ 4.2 Experimental Results
[…]
Ablation Study. Figure 3 shows three metrics in RL training. We compare our RLKD with GRPO because RLKD is actually GRPO with the reward from our Generative Structure Reward Model (GSRM). The results indicate that as the training progresses, GSRM enables RLKD to better optimize the accuracy of the task (Figure 3 (a)), primarily because the student gradually learns the teacher’s implicit multi-branch structure in reasoning (Figure 3 (c)). Please see Appendix for more ablation studies about hyparameters such as reward weights.
(a) Combined Accuracy (result acc. + format acc.)
@@
4.3 Analysis of RL distillation vs. SFT distillation
Performance Trend Varying with Training Steps.
@@
This section compares the SFT-based distillation with our RL-based distillation (RLKD) by training Qwen2.5-Math-7B on the dataset has domain shift to AIME24 and is out of the domain of GPQA. This setting, where there exists a domain discrepancy between the training and testing sets for distillation, allows us to intuitively discern whether SFT-based distillation is merely mimicking and memorizing the teacher’s paths, and whether our RL-based distillation enables the student to learn authentic reasoning. The experimental results are shown in Figure 4. It is noteworthy that, as the training progresses, SFT and RLKD demonstrate completely opposite performance trends: RLKD can consistently enhance performance, even when distilling on dataset that significantly diverges from the testing set, whereas SFT progressively undermines performance. It indicates that SFT distillation is easy to fall into the trap of simply imitating and memorizing the teacher’s reasoning paths, rather than learning authentic reasoning that can be stably generalized. Our RLKD makes the student learn to perform better sampling among potential branches at each step, which is the implicit multi-branch structure of teacher’s authentic reasoning.
@@
Figure 5: Diversity among different reasoning paths at each meta-reasoning step varying with training.
@@
The diversity of reasoning paths can reflect whether the student LLM has truly learned reasoning or is just memorizing the fixed teacher’s path in the training set. Figure 5 shows that our method brings the diversity-patterns of the student LLM closer to that of the teacher LLM in multi-step reasoning, suggesting that it allows the student LLM to learn authentic reasoning by distilling the implicit multi-branch structure. Specifically, we randomly sample 500 samples from Openthoughts-114K. On this sampled set, we set temperature to 0.80.80.8 and let LLM generate 161616 responses for each question. We use the GSRM in Section 3.1.1 to generate a corresponding sequence containing multiple meta-reasoning-solving steps for each response, and calculate the diversity of the meta-reasoning-solving content within the 161616 responses of each question at step-level. The diversity score $D$ is calculated as:
@@
in which $\mathbf{e}_{i}$ is the text embedding encoded by gte-Qwen2-7B-instruct [^5] https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct for each meta-reasoning-solving content. As shown in Figure 5, the teacher LLM (Deepseek-R1) has a significantly different diversity-pattern from the SFT-distilled student LLM (DeepSeek-R1-Distill-Qwen-7B). Reasoning in the teacher has the higher diversity while the student is stuck in the relatively fixed paths. As the training of our method progresses (from step 20 to 100), the diversity of the student’s reasoning paths begins to increase and gradually approaches that of the teacher, which indicates that our method allows the student to successfully learn the teacher’s authentic reasoning paradigm.
@@ 5 Conclusion and Discussion
Conclusion. This work addresses a critical flaw in knowledge distillation in LLM’s reasoning: the failure of SFT to transfer the implicit multi-branch structure underlying authentic reasoning. Drawing from cognitive neuroscience principles, we show that authentic reasoning involves dynamic meta-reasoning (sub-problem selection) and solving steps—a implicit multi-branch structure flattened by token-level SFT training. Our RLKD, the first RL-based distillation framework for reasoning, overcomes this paired with a Generative Structure Reward Model (GSRM), which decomposes reasoning paths into meta-reasoning-solving pairs and scores the structural alignment between teacher and student. Experiments across math and graduate-level QA tasks demonstrate RLKD’s superiority than SFT-based distillation, SFT-RL pipeline and RL baselines including PPO and GRPO, proving its ability to distill how teachers navigate latent reasoning branches rather than mimicking surface tokens. Further analysis confirms RLKD-trained students mirror teachers’ multi-branch exploration patterns, closing the imitation-authentic reasoning gap.
Limitations and Potential Impact. A key limitation of our work stems from computational constraints: while RLKD demonstrates effectiveness on 7B-scale models, its viability on larger LLMs (e.g., 70B+ parameters) remains unverified. The framework’s emphasis on structural distillation opens avenues for developing lightweight, generalizable LLMs, potentially reducing reliance on brute-force scaling—a timely direction given growing concerns about AI’s environmental and economic costs.
@@ Appendix B Ablation study for Reward Weights
In this section, we conduct ablation study about the reward weights in RL training. The essential rewards in our method include accuracy reward ($R_{acc}$) for the specific tasks such as rule-based accuracy reward in math, and our structure matching reward ($R_{gsrm}$) obtained from generative structure reward model, so we primarily discuss the weight relationship between these two rewards. Specifically, we adjust the weight between $R_{gsrm}$ and $R_{acc}$ ($\frac{R_{gsrm}}{R_{acc}}$) from 0.20.20.2 to 222 in increments of 0.20.20.2, and observe the variations in pass@1 (n=16) and pass@16 (n=16) on AIME24. This experimental results are shown in Figure 7. We can see that when the weights of the two are equal, i.e., $\frac{R_{gsrm}}{R_{acc}}=1$, the LLM achieves a harmonized optimal effect in both the correctness of reasoning (pass@1)and the capability of sampling (pass@16). So in our experiments, we set the reward weights as: $R_{acc}:R_{gsrm}:R_{format}:R_{tag}=3:3:2:2$, in which $R_{format}$ is format reward and $R_{tag}$ is tag count reward that assesses whether the thought content adheres to the specified format Guo et al. [2025].

### v4

@@ 3.1.1 Structured Reward Mechanism
After above generation, we denote $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$ as the sequences generated from reasoning paths of the teacher and the student LLM respectively:
[…]
We propose the Structured Reward Mechanism to map the generated sequence consisting of multiple meta-reasoning and solving to a reward value according to the alignment between $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$. Although this computation is carried out based on a linear sequence, when combined with RL, it becomes capable of quantifying the degree of alignment between the teacher and the student in terms of the implicit multi-branch structure, serving as environmental feedback in RL. This is because it can assess the step-level alignment between the sequences and, through the mechanism of reward for RL, finely guide the student LLM to make correct sampling from multiple potential sub-problems at each meta-reasoning step. Compared with SFT-based distillation, it makes the student LLM learn the implicit multi-branch structure in the authentic reasoning of the teacher LLM rather than just focuse on memorizing the reasoning path on the surface generated by the teacher LLM. Compared with existing RL methods, it can guide the student LLM on how to do better sampling at each step and avoid reward hacking Amodei et al. (2016); Di Langosco et al. (2022) by step-to-step comparison and early-exist mechanism. Specifically, our structured reward mechanism sequentially compares the corresponding steps of $\mathbf{S^{t}}$ and $\mathbf{S^{s}}$ (Algorithm 1). We use Qwen-2.5-7B-Instruct to determine whether two texts are matched. For the steps where $\mathcal{M}{{}^{t}_{i}}$ matches $\mathcal{M}{{}^{s}_{i}}$, we assign a temporary reward value of $1$. We further deduct the temporary reward by judging the matching relationship of $\mathcal{Q}_{i}$ and $\mathcal{A}_{i}$. The temporary reward value is added to the total reward value after this. When meeting the mismatched $\mathcal{M}_{i}$, the reward accumulation ends. This reward mechanism follows the sequential dependency of the reasoning path and can give an approximately unique reward value for each step.
@@ 3.2 RL-based Knowledge Distillation Training
We combine our Generative Structure Reward Models (GSRM) with Group-based Relative Policy Optimization (GRPO) Shao et al. (2024) for RL-based knowledge distillation training. In training, we combine the reward obtained from GSRM and the outcome reward of the specific task, such as the accuracy in math, in a weighted manner as the total reward for GRPO.
@@ 4 Experiments
We use OpenR1-Math as the training dataset for RL, in which Deepseek-R1 generated responses in this dataset are used as the teacher LLM’s reasoning paths. We keep the training datasets consistent with baselines including PPO and GRPO. In evaluation, we use the popular and challenging datasets on LLM’s reasoning including AIME24 MAA and MATH-500 Hendrycks et al. (2021) for math reasoning and GPQA-Diamond Rein et al. (2024) for graduate-Level Q&A. As for the metrics, we follow existing studies on LLM’s reasoning Guo et al. (2025); Shao et al. (2024); Zhang et al. (2025); Deng et al. (2024); Xu et al. (2024b) to use pass@k Chen et al. (2021). Our validation is divided into two parts: pass@1 and pass@k ($k>1$). As for pass@1, we generate $m$ ($m$ is $64$ for AIME, $8$ for GPQA and $4$ for MATH500) responses for each question and compute pass@1 as $pass@1=\frac{1}{m}\sum_{i=1}^{m}p_{i}$, in which $p_{i}$ is the correctness of the $i$-th response. This can alleviate randomness on small datasets. As for pass@k ($k>1$), it involves the LLM generating $k$ responses for each question, with the data sample being marked as accurate if at least one of the $k$ responses is accurate. Paying attention to the metric where $k>1$ is crucial, as in this setting, the LLM has the opportunity to explore multiple diverse paths to answer the question, it reflects the LLM’s ability to sample from multiple implicit paths during reasoning, thereby assessing whether the distilled LLM has learned authentic reasoning or merely memorized the teacher’s paths.
@@
We categorize the baselines into three groups according to the experimental settings. The first setting aims to show the effect of our method on improving LLM’s reasoning ability. We compare our RL-only training method with the LLM trained in SFT-RL pipeline. We use Qwen2.5-Math-7B-Instruct Yang et al. (2024), a powerful reasoning LLM trained on large-scale chain-of-thought datasets with SFT and then GRPO. Both our method and Qwen2.5-Math-7B-Instruct use the same based model: Qwen2.5-Math-7B, which has been pre-trained on math corpus. The second setting aims to explore whether our RL-based distillation can further improve the performance of the SFT-distilled LLM and achieve effective optimization of the SFT-RL pipeline in knowledge distillation. We compare our method with RL baselines including PPO and GRPO based on DeepSeek-R1-Distill-Qwen-7B Guo et al. (2025), a powerful LLM that is SFT-distilled from Deepseek-R1. The third setting aims to compare which method is better for the student LLM to learn authentic reasoning rather than memorizing the teacher’s path. We use embedding111obtained by gte-Qwen2-7B-instruct similarity Deng et al. (2025) to select the 3.2K data samples with the largest difference from the test set in OpenR1-Math-220k training set, and train Qwen2.5-Math-7B on this subset with RL-based distillation and SFT-based distillation respectively.
@@
In training, we build our code based on Open-R1, an open-source project for LLM’s reasoning. We use Pytorch 2.5.1 as the training framework and Deepspeed 0.15.4 for acceleration of parallel computing. In RL training, we use one Ascend 910B 64G NPU for online inference and four Ascend 910B 64G NPUs for training under deepspeed-zero3 setting. As for hyperparameters, we set per-device batch size as $2$, gradient accumulation steps as $4$, group size for GRPO as $4$, temperature in online inference as $0.7$. In evaluation, we use Lighteval as the toolkit and follow settings in Deepseek Guo et al. (2025) to set temperature as $0.6$, max new tokens as $32768$ and top-p as $0.95$. We report the results of multiple runs to reduce the randomness.
@@
(a) Performance on AIME24
@@
(b) Performance on GPQA-Diamond and MATH-500
means training set of MATH-500 has appeared in the SFT training data of Qwen2.5-Math-7B-Instruct Yang et al. (2024).
@@ 4.2 Experimental Results
Main Results. Results about reasoning abilities of LLMs are shown in Table 1. In the training based on Qwen2.5-Math-7B, our RL-only method Qwen2.5-Math-7B-RLKD-Zero outperforms complex SFT + RL pipeline (Qwen2.5-Math-7B-Instruct) and uses much less data (nearly $0.1\%$). In the training based on Deepseek-R1-Distill-Qwen-7B (SFT-distilled LLM), baseline RL methods including PPO and GRPO can hardly bring about significant improvements on this basis while our RLKD is capable of further enhancing performance. The indicates that our RL-based distillation approach enables the SFT-distilled LLM to learn additional information beyond its memorization of the teacher’s reasoning paths. The relatively more significant improvements are observed in pass@k ($k>1$). In this setting, LLM has the opportunity to explore multiple diverse paths to answer the question, which means that compared to SFT distillation, our method enables the student LLM to learn how to perform sampling from multiple potential paths by distilling the implicit multi-branch structure from the teacher, thereby increasing the probability of providing the correct answer.
Ablation Study. Figure 3 shows three metrics in RL training. We compare our RLKD with GRPO because RLKD is actually GRPO with the reward from our Generative Structure Reward Model (GSRM). The results indicate that as the training progresses, GSRM enables RLKD to better optimize the accuracy of the task (Figure 3 (a)), primarily because the student gradually learns the teacher’s implicit multi-branch structure in reasoning (Figure 3 (c)).
(a) Accuracy
@@
Figure 5: Diversity among different reasoning paths at each meta-reasoning step varying with training. Red line is Deepseek-R1. Blue line is DeepSeek-R1-Distill-Qwen-7B. Green line is DeepSeek-R1-Distill-Qwen-7B w/ RLKD.
4.3 Compare RL Distillation with SFT Distillation
Performance Trend Varying with Training Steps.
This section compares the SFT-based distillation with our RL-based distillation (RLKD) by training Qwen2.5-Math-7B on the dataset has domain shift to AIME24 and is out of the domain of GPQA. This setting, where there exists a domain discrepancy between the training and testing sets for distillation, allows us to intuitively discern whether SFT-based distillation is merely mimicking and memorizing the teacher’s paths, and whether our RL-based distillation enables the student to learn authentic reasoning. The experimental results are shown in Figure 4. It is noteworthy that, as the training progresses, SFT and RLKD demonstrate completely opposite performance trends: RLKD can consistently enhance performance, even when distilling on dataset that significantly diverges from the testing set, whereas SFT progressively undermines performance. It indicates that SFT distillation is easy to fall into the trap of simply imitating and memorizing the teacher’s reasoning paths, rather than learning authentic reasoning that can be stably generalized. Our RLKD teaches students to sample from multiple potential branches, much like the multi-branch structure implicit in real reasoning in teacher LLMs.
@@ 4.3 Compare RL Distillation with SFT Distillation
The diversity of reasoning paths can reflect whether the student LLM has truly learned reasoning or is just memorizing the fixed teacher’s path in the training set. Figure 5 shows that our method brings the diversity-patterns of the student LLM closer to that of the teacher LLM in multi-step reasoning, suggesting that it allows the student LLM to learn authentic reasoning by distilling the implicit multi-branch structure. Specifically, we randomly sample 500 samples from Openthoughts-114K. On this sampled set, we set temperature to $0.8$ and let LLM generate $16$ responses for each question. We use the GSRM in Section 3.1 to generate a corresponding sequence containing multiple meta-reasoning-solving steps for each response, and calculate the diversity of the meta-reasoning-solving content within the $16$ responses of each question at step-level. The diversity score $D$ is calculated as:
@@
in which $\mathbf{e}_{i}$ is the text embedding encoded by gte-Qwen2-7B-instruct for each meta-reasoning-solving content. As shown in Figure 5, the teacher LLM (Deepseek-R1) has a significantly different diversity-pattern from the SFT-distilled student LLM (DeepSeek-R1-Distill-Qwen-7B). Reasoning in the teacher has the higher diversity while the student is stuck in the relatively fixed paths. As the training of our method progresses (from step 20 to 100), the diversity of the student’s reasoning paths begins to increase and gradually approaches that of the teacher, which indicates that our method allows the student to learn the teacher’s authentic reasoning paradigm.
@@ 5 Conclusion and Discussion
This work addresses a critical flaw in knowledge distillation in LLM’s reasoning: the failure of SFT to transfer the implicit multi-branch structure underlying authentic reasoning. Drawing from cognitive neuroscience principles, we show that authentic reasoning involves dynamic meta-reasoning (sub-problem selection) and solving steps—a implicit multi-branch structure flattened by token-level SFT training. Our RLKD, the first RL-based distillation framework for reasoning, overcomes this paired with a Generative Structure Reward Model (GSRM), which decomposes reasoning paths into meta-reasoning-solving pairs and scores the structural alignment between teacher and student. Experiments across math and graduate-level QA tasks demonstrate RLKD’s superiority than SFT-based distillation, SFT-RL pipeline and RL baselines including PPO and GRPO, proving its ability to distill how teachers navigate latent reasoning branches rather than mimicking surface tokens. Further analysis confirms RLKD-trained students mirror teachers’ multi-branch exploration patterns, closing the imitation-authentic reasoning gap.
Acknowledgments
This work was supported by the Key Research and Development Program of Xinjiang Uyghur Autonomous Region Grant No. 2024B03026, the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the Beijing Nova Program under Grants No. 20250484765, the National Natural Science Foundation of China (NSFC) under Grants No. 62276248, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.
@@ Appendix B Ablation study for Reward Weights
In this section, we conduct ablation study about the reward weights in RL training. The essential rewards in our method include accuracy reward ($R_{acc}$) for the specific tasks such as rule-based accuracy reward in math, and our structure matching reward ($R_{gsrm}$) obtained from generative structure reward model, so we primarily discuss the weight relationship between these two rewards. Specifically, we adjust the weight between $R_{gsrm}$ and $R_{acc}$ ($\frac{R_{gsrm}}{R_{acc}}$) from $0.2$ to $2$ in increments of $0.2$, and observe the variations in pass@1 (n=16) and pass@16 (n=16) on AIME24. This experimental results are shown in Figure 7. We can see that when the weights of the two are equal, i.e., $\frac{R_{gsrm}}{R_{acc}}=1$, the LLM achieves a harmonized optimal effect in both the correctness of reasoning (pass@1)and the capability of sampling (pass@16). So in our experiments, we set the reward weights as: $R_{acc}:R_{gsrm}:R_{format}:R_{tag}=3:3:2:2$, in which $R_{format}$ is format reward and $R_{tag}$ is tag count reward that assesses whether the thought content adheres to the specified format Guo et al. [2025].